wrkwhrkhr:

Solve the following system of equations and show all work. y = x2 + 3 y = x + 5

CripQUEZZ:

is it supposed to be x squared? or 2x

GalaxyzStarz:

\[y=x^2+3\] \[y=x+5\]

GalaxyzStarz:

Since they both equal y, set both of them equal to each other. \[x^2+3=x+5\] \[x^2-x-2=0\] \[(x-2)(x+1)=0\] The two solutions would be: \[x-2=0\] \[x=2\] and \[x+1=0\] \[x=-1\] I hope this cleared up a bit of confusion.

SwaggyMark:

both equal y meaning that the equations are equal to each other so x^2 + 3 = x + 5 because there is an exponent we are going to have to put the equation in quadratic form ax^2 + + bx + c x^2 + 3 = x + 5 subtract x from both sides x^2 - x + 3 = 5 subtract 5 from both sides x^2 - x - 2 we can then factor to get (x+1)(x-2)=0 we then solve each factor x-2 = 0 x + 1 = 0 add 2 to both sides subtract 1 from both sides x = 2 and x = -1

GalaxyzStarz:

@wrkwhrkhr wrote:
@galaxyzstarz ,it did actually,and thank you overall.You did in fact deserved that medal.As for @cripquezz ,you deserve a medal too for still trying to help;I present you an imaginary medal.
He'll/She'll earn my medal.

wrkwhrkhr:

As well as to @swaggymark ,sorry about that;and to everyone else that tried their best to help.Farewell to all.

CripQUEZZ:

@wrkwhrkhr wrote:
@galaxyzstarz ,it did actually,and thank you overall.You did in fact deserved that medal.As for @cripquezz ,you deserve a medal too for still trying to help;I present you an imaginary medal.
lol i was going to help but had to do sum but thank you for the medal